OpenStudy (anonymous):

Could you explain restrictions to the domain and range of a sequence.(arithmetic and geometric sequence) please...

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

been stuck on this for a while. :(

OpenStudy (perl):

For me it is easiest to think of sequences as functions with domain of integers. for example the sequence f(n) = 2*n, lets assume the domain is n=1,2,3,.. that gives us the sequence 2*1, 2*2, 2*3, 2*4, ... 2*n , ...

OpenStudy (anonymous):

okay, so here are my numbers 3,6,9 and 4,8,16

OpenStudy (anonymous):

still there @perl?

OpenStudy (perl):

yes

OpenStudy (perl):

3,6,9 can be represented by the rule f(n) = 3*n the domain will be 1,2,3 in that case

OpenStudy (anonymous):

have any ideas? i'm very lost...

OpenStudy (anonymous):

oooohhhh, i get it thank u soooo much @perl :)